Please adhere to the rounding instructions(in red) for calculations and final answer. If the answer is not exact, it will be incorrect. Thanks!

Question 6 2 pts Investment X offers to pay you $2910 per year for 15 years, whereas Investment Y offers to pay you $8660 per year for 3 years. Both investments have a discount rate of 12.3%. Calculate the present value for Investment Y. (Round answer to 2 decimal places. Do not round intermediate calculations) Topic: Discounted Cash Flow Valuation
Answer #1

Present value factor

= 1 / (1 + r) ^ n

Where,

r = Rate of interest = 12.3% or 0.123

n = Years = 0 to 3

So, Present value factor for year 2 will be

= 1 / (1.123) ^ 2

= 1 / 1.261129

= 0.792940

Following table shows the calculation of present value of investment Y

Calculations Particulars
Years 0 1 2 3
A Cash Flows 0 8660 8660 8660
B Present value factor 1 0.890472 0.79294 0.706091
C = A x B Present value 0 7711.487 6866.863 6114.749
D = Sum C Total present value 20693.10

So, the present value of investment Y is $ 20,693.10